# Verifying the failing supernova constraint on dark photons with two-dimensional hydrodynamic simulations

### v1

@@ IV SUMMARY and Discussion
In this study, we performed two-dimensional neutrino-radiation hydrodynamic simulations for a collapsing star coupled with the production of DPs with $m_{\gamma^{\prime}}=0.3$ MeV. As the photon-DP mixing parameter is increased, the successful supernova explosion transitioned to failures in the explosion. This suggests that the mixing parameter can be constrained by the very existence of supernova explosions in the Universe, as proposed by 2025PhRvL.134o1002C. Our models show quantitative agreement with the constraint obtained in 2025PhRvL.134o1002C in the case of $m_{\gamma^{\prime}}=0.3$ MeV, encouraging to extend the argument to various exotic particle models.
@@
Supernova explodability is dependent on the spatial dimension of simulations. A way to quantify explodability is the critical neutrino luminosity as a function of the mass accretion rate, above which steady state solutions do not exist [1993ApJ...416L..75B, 2005ApJ...623.1000Y]. Multi-dimensional hydrodynamic simulations have revealed that the critical luminosity is the highest for one-dimensional models and the lowest for two-dimensional models, and the one for three-dimensional models is between the two [2008ApJ...688.1159M, 2012ApJ...755..138H, 2012ApJ...749...98T, 2013ApJ...775...35C]. This suggests that three-dimensional models may explode even if we consider higher values of the photon-DP mixing parameter. In this study, we developed two-dimensional models to reduce the computational cost. Our constraint is conservative in a sense that the models are easier to explode than three-dimensional models. It is still desirable to perform three-dimensional simulations coupled with DPs in the future to obtain a more stringent constraint.
@@
It is also important to extend this study to investigate a wider DP mass range. In our simulations, the DP mass is fixed to $m_{\gamma^{\prime}}=0.3$ MeV. However, the failing supernova constraint is stronger than the cooling constraint for DPs with $m_{\gamma^{\prime}}\sim 0.1$–0.4 MeV [2025PhRvL.134o1002C]. Although our results showed that simulations and post-processing lead to a similar result when $m_{\gamma^{\prime}}=0.3$ MeV, it is not guaranteed that simulations can reproduce the previous result for other DP masses. In addition, DPs heavier than $\sim 1$ MeV can decay into a electron-positron pair. In this case, DPs can contribute to the energy transfer in a star and enhance the explosion energy. In our simulations, we did not consider the DP heating rate because the DP mean free path is long enough. However, when DPs are unstable or the mixing parameter is much higher, hydrodynamic simulations that solve the DP transfer consistently should be performed, as done for other exotic particles such as axions [2022JCAP...08..045C, 2022PhRvD.105f3009M, 2023PhRvD.108f3027M, 2025PhRvD.111j3028T] and sterile neutrinos [2014PhRvD..90j3007W, 2018PhRvD..98j3010R, 2024PhRvD.110b3031M].
@@
Last but not least, it is worthwhile to perform simulations coupled with DPs for a wide range of progenitors to investigate the DP effect on the mass function of neutron stars and black holes. Surveying the vast parameter space with multi-dimensional simulations will require a large amount of computational resources. To circumvent this problem, one may perform the so-called 1D+ simulations, which take turbulence into account phenomenologically [2019ApJ...887...43M, 2019MNRAS.487.5304M, 2020ApJ...890..127C, 2021ApJ...912...29B, 2022ApJ...926..147B, 2024MNRAS.528.1158S].
@@ Acknowledgements.
Numerical computations were carried out on Cray XD2000 at Center for Computational Astrophysics, National Astronomical Observatory of Japan. This work was in part supported by JSPS KAKENHI Grant Numbers K.M.: JP23KJ2147, JP23K13107, JP25H02194; $T.$: JP23K25895, JP23K22494, JP23K03400, JP24K00631; $K.$: JP23KF0289, JP24K07027 and MEXT KAKENHI Grants Number JP24H01825 ($K.$). This work was also supported by “Program for Promoting researches on the Supercomputer Fugaku” (Structure and Evolution of the Universe Unraveled by Fusion of Simulation and AI; Grant Number JPMXP1020230406), and JICFuS.

### v2

@@ IV SUMMARY and Discussion
In this study, we performed two-dimensional neutrino-radiation hydrodynamic simulations for a collapsing star coupled with the production of DPs with $m_{\gamma^{\prime}}=0.3$ and 0.45 MeV. As the photon-DP mixing parameter is increased, the successful supernova explosion transitioned to failures in the explosion. This suggests that the mixing parameter can be constrained by the very existence of supernova explosions in the Universe, as proposed by Caputo et al. [21, C25]. Our models show quantitative agreement with the constraint obtained in C25 in the case of $m_{\gamma^{\prime}}=0.3$ MeV, encouraging to extend the argument to various exotic particle models.
@@
Supernova explodability is dependent on the spatial dimension of simulations. A way to quantify explodability is the critical neutrino luminosity as a function of the mass accretion rate, above which steady state solutions do not exist [16, 91]. Multi-dimensional hydrodynamic simulations have revealed that the critical luminosity is the highest for one-dimensional models and the lowest for two-dimensional models, and the one for three-dimensional models is between the two [57, 37, 83, 27]. This suggests that three-dimensional models may explode even if we consider higher values of the photon-DP mixing parameter. In this study, we developed two-dimensional models to reduce the computational cost. Our constraint is conservative in a sense that the models are easier to explode than three-dimensional models. It is still desirable to perform three-dimensional simulations coupled with DPs in the future, because they are expected to lead to a more stringent constraint.
@@
It is also important to extend this study to investigate a wider DP mass range. In our simulations, the DP mass is fixed to $m_{\gamma^{\prime}}=0.3$ and 0.45 MeV. However, the failing supernova constraint is stronger than the cooling constraint for DPs with $m_{\gamma^{\prime}}\sim 0.1$–0.4 MeV [21]. Although our results showed that simulations and post-processing lead to a similar result when $m_{\gamma^{\prime}}=0.3$ MeV, it is not guaranteed that simulations can reproduce the previous result for other DP masses. In fact, our simulations suggest that, when $m_{\gamma^{\prime}}=0.45$ MeV, the mixing parameter should satisfy $\epsilon<3\times 10^{-9}$ to obtain successful supernova explosions, while the parameter is located in the allowed region based on the post-processing. In addition, DPs heavier than $\sim 1$ MeV can decay into a electron-positron pair. In this case, DPs can contribute to the energy transfer in a star and enhance the explosion energy. In our simulations, we did not consider the DP heating rate because the DP mean free path is long enough. However, when DPs are unstable or the mixing parameter is much higher, hydrodynamic simulations that solve the DP transfer consistently should be performed, as done for other exotic particles such as axions [24, 52, 53, 82] and sterile neutrinos [89, 70, 54].
@@
Last but not least, it is worthwhile to perform simulations coupled with DPs for a wide range of progenitors to investigate the DP effect on the mass function of neutron stars and black holes. Surveying the vast parameter space with multi-dimensional simulations will require a large amount of computational resources. To circumvent this problem, one may perform the so-called 1D+ simulations, which take turbulence into account phenomenologically [47, 56, 26, 11, 12, 73].
@@ Acknowledgements.
Numerical computations were carried out on Cray XD2000 at Center for Computational Astrophysics, National Astronomical Observatory of Japan. This work was in part supported by JSPS KAKENHI Grant Numbers K.M.: JP23KJ2147, JP23K13107, JP25H02194; T.T.: JP23K25895, JP23K22494, JP23K03400, JP24K00631; K.K.: JP23KF0289, JP24K07027 and MEXT KAKENHI Grants Number JP24H01825 (K.K.). This work was also supported by “Program for Promoting researches on the Supercomputer Fugaku” (Structure and Evolution of the Universe Unraveled by Fusion of Simulation and AI; Grant Number JPMXP1020230406), and JICFuS.